# Spin waves in Co2FeGe films

### v1

@@ 2.2 Measurements
Brillouin Light Scattering (BLS) measurements were performed at RT using a six-pass tandem Brillouin spectrometer (TFP2-HC, JRS) [25, 28, 29, 30]. The light was generated by a frequency-stabilized DPSS laser operating at $\lambda_{0}$ = 532 nm (Coherent Verdi V5). Measurements were made in the backscattering geometry with the light polarization ps for magnons. The backscattered light was collected using f/8 optics with a focal length of 150 mm. The external DC field ($H$) was applied in the sample plane and perpendicular to the plane of light incidence, i.e., perpendicular to the wave vector k. This geometry is called Damon-Eshbach or magnetostatic surface wave geometry; Fig. 1. Angle $\theta$ is the angle between the incident light and the normal to the sample plane, and the scattering wave vector k was defined by the wavelength of the incident light and $\theta$ as $k=\frac{4\pi}{\lambda_{0}}\sin{\theta}$. The wave vector range was $0.01\div 2.25\times 10^{7}\ \mathrm{m}^{-1}$, with a resolution of $0.01\times 10^{7}\ \mathrm{m}^{-1}$. Each spectrum was accumulated for 1 hour, and each peak in the spectrum was fitted using a Lorentzian curve. The frequency resolution in the BLS experiment was about 0.07 GHz.
Broadband ferromagnetic resonance (FMR) measurements were performed at RT utilizing a coplanar waveguide connected to a vector network analyzer operating in the frequency ($f$) range 50 MHz - 20 GHz. The external DC field ($H$) was applied in the plane of the films. The microwave field excitation ($\sim 0.1$ Oe) was applied in-plane and perpendicular to the DC field. The frequency spectra of the complex magnetic susceptibility $\chi(f)$ were subsequently derived from the raw S21 data using the practices outlined in references [31, 32].
@@ 3.1 Structure
The structure of the films under investigation was intensively studied by X-ray diffraction techniques in our previous work [24] . Out-of-plane (001) Co2FeGe —— (001) MgO and in-plane [110] Co2FeGe —— [100] MgO epitaxial growth conditions were confirmed. Deposition at the elevated temperature and heat treatment cause improvements of crystal quality and atomic ordering. For sample RT B2-type atomic ordering was observed, with Co atoms occupying their respective lattice positions while Fe and Ge were intermixed. Superlattice all odd (hkl) index peaks (111) and (113), characteristic of fully atomically ordered L21 phase, appeared in X-ray diffractograms for samples HT and RT-A. However, these reflections are also present for partially ordered DO3, phase (Co and Fe or Co and Ge atoms are mixed at their crystallographic positions). Distinguishing between the L21 and DO3 phases using conventional X-ray techniques is challenging due to the nearly identical atomic scattering factors of Co, Fe, and Ge atoms. [33]. Although the formation of L21 atomic ordering is desirable due to its predicted half-metallicity [9], it is not always consistently achieved during fabrication [34]. Sample RT exhibits a slight tetragonal distortion of Co2FeGe cubic structure caused by in-plane tensile strain due to the lattice mismatch between Co2FeGe and MgO. The diagonal of the MgO unit cell is  0.5958 nm, while the bulk lattice parameter of Co2FeGe is 0.5738 nm. As a result, the in-plane lattice parameter of the film (a = 0.5759 nm) was found to be larger, while the out-of-plane parameter (c = 0.5671 nm) was smaller compared to the bulk, with a c/a ratio of  0.985. This distortion is reduced for samples HT and RT-A. The unit cell becomes almost cubic, manifesting the relaxation of the strain (c/a 0.999 for sample HT and  0.99 for sample RT-A). Analysis of the X-ray rocking curve widths suggests that the epitaxial quality of the films improves on heating the substrate or annealing. Furthermore, X-ray reflectivity measurements show a reduction in surface roughness (from 1.4 nm for sample RT to  0.9 nm for samples HT and RT-A) and an increase in film density (from 8.26 g/cm3 to 8.3 g/cm3, respectively). All of these factors point to the improvement of the microstructure of the films produced by heat treatment.
@@ 3.2 Brillouin light scattering
Fitting the combined set of the $\omega(H,k)$ data shown in Figs. 3,4 using the Kalinikos-Slavin formalism [39] yields higher values for the effective magnetization $M$ and the exchange stiffness constant $A$ for samples HT and RT-A compared to sample RT, with the values shown in the respective panels of Fig. 3. This indicates that deposition on a heated substrate or deposition at room temperature with subsequent annealing improves the crystal structure and, thereby, increases the degree of atomic and magnetic ordering in the material. Specifically, the obtained significantly enhanced $M$ and $A$ for sample RT-A associate with the observed higher degree of structural ordering in the L21 phase of the material (see Structure), known to have highly spin-polarized conductivity (approaching semi-metallic), which is attractive for spintronic devices such as sensors and oscillators.
Both surfaces of our films are exposed to oxygen, either in the MgO substrate (bottom surface) or air (top surface). As an example, the interaction between CoFeB and MgO in magnetic tunnel junctions [38] is well known to produce out-of-plane magnetic anisotropy used in spin-transfer-torque devices such as MRAM. Another common origin of surface anisotropy is interfacial strains [40], seen in our samples as detailed in sec. Structure, asymmetric at bottom vs. top surface. With multiple anisotropy contributions typically present in a ferromagnetic film, it is often non-trivial to extract just the surface components, especially to differentiate them. Our BLS data were analyzed using the Kalinikos-Slavin theory [39]. Its free-surface-spins case is found to fit the experimental data somewhat better than the pinned-surface-spins case (both cases symmetric regarding top vs. bottom), although both yield good fits to the experimental data for $\omega(H)$ of Fig. 3 as well as the DE mode in $\omega(k)$ of Fig. 4. The corresponding fits for the PSSW modes in $\omega(k)$ were not satisfactory, which is attributed to likely issues with crossing between different SW-modes [41]. Analysis of such cases is non-trivial and goes beyond the scope of this paper. The overall quality of the fitting procedure used, as it relates to extracting reliable material parameters, is nevertheless judged to be good, and is verified with a complementary FMR technique (see FMR section below).
@@
$\mathrm{FWHM}=2\pi\alpha\gamma H+\delta f_{0},\$ (1)
where $\alpha$ represents the intrinsic spin damping (Gilbert damping parameter), and $\delta f_{0}$ accounts for extrinsic contributions to the linewidth, unrelated to the magnetic field, primarily due to instrumentational limitations and sample inhomogeneity. The obtained $\alpha$ values for as deposited (RT) and annealed (RT-A) samples are shown in Fig. 5 and give additional evidence for improved material quality (lower damping) after thermal treatment. Linear fitting yielded the correlation coefficients of 0.99 (RT) and 0.97 (RT-A), indicating high fit accuracy (no reliable fit could be obtained for sample HT, however). Thus, a significant reduction in the magnetization dissipation is clearly observed on annealing, which is another evidence of improved microstructure. The $\alpha$ values obtained and the trend on annealing are in agreement with our resent measurements using cavity-FMR [24].

### v2

@@ 2.2 Measurements
Brillouin Light Scattering (BLS) is an effective tool to characterize magnetization dynamics in Heusler alloys, recently used to study spin-waves in thick (120 nm) films of Co2Mn0.6Fe0.4Si [28]. Here, measurements were performed at RT using a six-pass tandem Brillouin spectrometer (TFP2-HC, JRS) [25, 29, 30]. The light was generated by a frequency-stabilized DPSS laser operating at $\lambda_{0}$ = 532 nm (Coherent Verdi V5). Measurements were made in the backscattering geometry with the light polarization ps for magnons. The backscattered light was collected using f/8 optics with a focal length of 150 mm. The external DC field ($H$) was applied in the sample plane and perpendicular to the plane of light incidence, i.e., perpendicular to the wave vector k. This geometry is called Damon-Eshbach or magnetostatic surface wave geometry; Fig. 1. Angle $\theta$ is the angle between the incident light and the normal to the sample plane, and the scattering wave vector k was defined by the wavelength of the incident light and $\theta$ as $k=\frac{4\pi}{\lambda_{0}}\sin{\theta}$. The wave vector range was $0.01\div 2.25\times 10^{7}\ \mathrm{m}^{-1}$, with a resolution of $0.01\times 10^{7}\ \mathrm{m}^{-1}$. Each spectrum was accumulated for 1 hour, and each peak in the spectrum was fitted using a Lorentzian curve. The frequency resolution in the BLS experiment was about 0.07 GHz.
Broadband ferromagnetic resonance (FMR) measurements were performed at RT utilizing a coplanar waveguide connected to a vector network analyzer Anritsu 37247 (Anritsu Corp.) operating in the frequency range 50 MHz – 20 GHz. The external DC field ($H$) was applied in the plane of the films. The microwave field excitation ($\sim 0.1$ Oe) was applied in-plane and perpendicular to the DC field. The frequency spectra of the complex magnetic susceptibility $\chi(f)$ were subsequently derived from the raw S21 data using the practices outlined in references [31, 32].
@@ 3.1 Structure
The structure of the films under investigation was intensively studied by X-ray diffraction techniques in our previous work [24] . Out-of-plane (001) Co2FeGe ||||| | (001) MgO and in-plane [110] Co2FeGe ||||| | [100] MgO epitaxial growth conditions were confirmed. Deposition at the elevated temperature and heat treatment cause improvements of crystal quality and atomic ordering. For sample RT B2-type atomic ordering was observed, with Co atoms occupying their respective lattice positions while Fe and Ge were intermixed. Superlattice peaks with odd (hkl) indices (e.g., (111) and (113)), characteristic of fully atomically ordered L21 phase, appeared in X-ray diffractograms for samples HT and RT-A. However, these reflections are also present for partially ordered DO3, phase (Co and Fe or Co and Ge atoms are mixed at their crystallographic positions). Distinguishing between the L21 and DO3 phases using conventional X-ray techniques is challenging due to the nearly identical atomic scattering factors of Co, Fe, and Ge atoms. [33]. Although the formation of L21 atomic ordering is desirable due to its predicted half-metallicity [9], it is not always consistently achieved during fabrication [34]. Sample RT exhibits a slight tetragonal distortion of Co2FeGe cubic structure caused by in-plane tensile strain due to the lattice mismatch between Co2FeGe and MgO. The diagonal of the MgO unit cell is  0.5958 nm, while the bulk lattice parameter of Co2FeGe is 0.5738 nm. As a result, the in-plane lattice parameter of the film (a = 0.5759 nm) was found to be larger, while the out-of-plane parameter (c = 0.5671 nm) was smaller compared to the bulk, with a c/a ratio of  0.985. This distortion is reduced for samples HT and RT-A. The unit cell becomes almost cubic, manifesting the relaxation of the strain (c/a 0.999 for sample HT and  0.99 for sample RT-A). Analysis of the X-ray rocking curve widths suggests that the epitaxial quality of the films improves on heating the substrate or annealing. Furthermore, X-ray reflectivity measurements show a reduction in surface roughness (from 1.4 nm for sample RT to  0.9 nm for samples HT and RT-A) and an increase in film density (from 8.26 g/cm3 to 8.3 g/cm3, respectively). All of these factors point to the improvement of the microstructure of the films produced by heat treatment.
@@ 3.2 Brillouin light scattering
Fitting the combined set of the $\omega(H,k)$ data shown in Figs. 3,4 using the Kalinikos-Slavin formalism [39] yields higher values for the effective magnetization $M$ and the exchange stiffness constant $A$ for samples HT and RT-A compared to sample RT  [^1] We previously determined using cavity FMR measurements [24] that the samples under study possess fourth-order in-plane magnetocrystalline anisotropy: $H_{a}$(RT) = -55 Oe, $H_{a}$(HT)=27 Oe, and $H_{a}$(RT-A)=-35 Oe; these values were taken into account in fitting the FMR-BLS data, with the values shown in the respective panels of Fig. 3. This indicates that deposition on a heated substrate or deposition at room temperature with subsequent annealing improves the crystal structure and, thereby, increases the degree of atomic and magnetic ordering in the material. Specifically, the obtained significantly enhanced $M$ and $A$ for sample RT-A associate with the observed higher degree of structural ordering in the L21 phase of the material (see Structure), known to have highly spin-polarized conductivity (approaching semi-metallic), which is attractive for spintronic devices such as sensors and oscillators.
The free-surface-spins case of the Kalinikos-Slavin theory [39] is found to fit our BLS data somewhat better than the pinned-surface-spins case (both cases symmetric regarding top vs. bottom), although both yield good fits to the experimental data for $\omega(H)$ of Fig. 3 as well as the DE mode in $\omega(k)$ of Fig. 4. The corresponding fits for the PSSW modes in $\omega(k)$ were not satisfactory, which is attributed to likely issues with crossing between different SW-modes [40]. Analysis of such cases is non-trivial and goes beyond the scope of this paper. The overall quality of the fitting procedure used, as it relates to extracting reliable material parameters, is nevertheless judged to be good, and is verified with a complementary FMR technique (see FMR section below).
@@
$\mathrm{FWHM}=\frac{2\alpha\gamma}{\pi}H+\delta f_{0},\$ (1)
where $\alpha$ represents the intrinsic spin damping (Gilbert damping parameter), and $\delta f_{0}$ accounts for extrinsic contributions to the linewidth, unrelated to the magnetic field, primarily due to instrumentational limitations and sample inhomogeneity. The obtained $\alpha$ values for as deposited (RT) and annealed (RT-A) samples are shown in Fig. 5 and give additional evidence for improved material quality (lower damping) after thermal treatment. Linear fitting yielded the correlation coefficients of 0.99 (RT) and 0.97 (RT-A), indicating high fit accuracy (no reliable fit could be obtained for sample HT, however). Thus, a significant reduction in the magnetization dissipation is clearly observed on annealing, which is another evidence of improved microstructure. The $\alpha$ values obtained and the trend on annealing obtained using BLS are in agreement with our recent measurements using broadband-FMR [24]: in excellent agreement for sample RT (0.01 vs. 0.089) and satisfactory agreement for sample RT-A (0.0066 vs. 0.0042); BLS had no satisfactory fit for sample HT whereas FMR yielded $\alpha=0.068$.
@@ 3.3 Ferromagnetic resonance
It may be informative to note that, from our broadband-FMR measurements, the inhomogeneous contribution to the resonance linewidth was estimated to be 62 Oe for the RT sample and 38 Oe for the RT-A sample. Such small values indicate that the films are essentially free from structural defects. Considering additionally their relatively large thickness, surface defects can be also neglected. Therefore, two-magnon scattering, caused predominantly by structural imperfections[47], is unlikely to contribute to the spin-damping values observed for the studied films.